# On the strong DR/DZ equivalence conjecture

### v1

@@ 2.5. The $A=B$ identity in the Gorenstein quotient
As a corollary, the first statement of this theorem implies the polynomiality of the system of Dubrovin-Zhang conservation laws associated to an arbitrary F-CohFT whose classes belong to the tautological ring. The second statement implies that the latter hierarchy is Miura equivalent, in a very explicit way, to the DR hierarchy of conservation laws associated to the same F-CohFT, also under assumption that its classes belong to the tautological ring. Finally, the third statement implies the strong DR/DZ correspondence for (partial) CohFTs, also under assumption that the classes of (partial) CohFT belong to the tautological ring. See [16, Section 4] and [6, 8].
@@ 2.6. The trivial CohFT
The second and the third statement are literally equivalent to the statements of the strong DR/DZ correspondence in this case, see [5, 6, 16]. The first statement is equivalent by [16, Section 4] to the fact that the second and higher derivatives of the potential $F(t_{0},t_{1},\dots)$ corresponding to the string solution are known to be differential polynomials in the dependent variables in this case, see e. g. [18]. ∎
@@ 3.1.2. Classes assigned to vertices
where the negative cohomological degree is just formally assigned to allow to treat this case non-exceptionally in what follows.
For a non-root root vertex $v$, we consider the class
@@
Note that in this case $a(v)=-a(h_{H(v)})=a(h_{1})+\cdots+a(h_{|H_{+}(v)|})$. In the exceptional unstable case $g(v)=0,H_{+}(v)=1$ we formally assign to this vertex the following class:
@@
where, as it was for the root vertex, the negative cohomological degree is just formally assigned to allow to treat this case non-exceptionally in what follows.
@@ 3.1.3. Classes assigned to trees
(3.5) $\Xi(T)\coloneqq u^{2g-2+m}\left(\prod_{e\in E(T)}\frac{a(e)}{u}\right)(b_{T})_{*}\left(\left(\sum_{d=-1}^{\infty}\frac{\Psi(v_{r})_{d}}{(-u)^{d}}\right)\otimes\bigotimes_{v\in V\setminus\{v_{r}\}}\left(\sum_{d=-1}^{\infty}\frac{{\mathbb{D}}(v)_{d}}{u^{d}}\right)\right),$
where $(b_{T})_{*}$ is the boundary pushforward map from $\bigotimes_{v\in V(T)}R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})$ to $R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ extended by linearity to the rational functions in $a_{i}$ and Laurent polynomials in $u$. In the case of pre-stable trees $b_{T}$ is assumed to contract the unstable components. Note that under the assumption $02g-2+n+m>$ 0 the dependence of the resulting formula on $a_{i}$’s is purely polynomial, so $\Xi(T)\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q[u,u^{-1}]$.
@@
Note that $[u^{2g-2+m-d}]\Xi(T)\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $\Xi(T)_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
@@ 3.1.4. Classes assigned to sets of trees
Note that $(\Xi^{m}_{g,n})_{d}\coloneqq[u^{2g-2+m-d}]\Xi^{m}_{g,n}\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $(\Xi^{m}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
@@ 3.2. Intersection with a divisor
Let $\rho_{2}\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+m_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+m_{2}+1}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map and assume that $m_{1}$ and $m_{2}$ are non-zero.
@@
Let $\rho_{3}\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+m+1}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map and assume that all points with the labels ${n+1},\dots,{n+m}$ are on the first component, and $n_{2}$ points on the second component have labels in $I\subset\{1,\dots,n\}$, $|I|=n_{2}$.
@@
$\quad+(\Xi^{m}_{g_{1},n_{1}+1})_{d-2g_{2}}(\{a_{j}\}_{j\not\in I},a_{I})\otimes\lambda_{g_{2}}\mathrm{DR}_{g_{2}}\Big{(}\{a_{i}\}_{i\in I},-a_{I}\Big{)}.$
In this formula we assume that the arguments of $\Xi$-classes correspond to the marked points on the corresponding components. In the last summand we use $a_{i}\coloneqq\sum\nolimits_{i\in I}a_{i}$.
@@ 3.3. Intersection with a kappa class
Lemmata in Section 3.2 reduce the computation of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha\Xi^{m}_{g,n}$ for an arbitrary class $\alpha$ to the computation of the same intersection numbers for the classes $\alpha=\kappa_{c_{1},\dots,c_{l}}\prod_{i=1}^{n+m}\psi_{i}^{b_{i}}$. To this end, we have the following statement:
@@ 3.4. Intersection with a $\psi$ class at a regular leg
where $a=\sum_{i=1}^{n}a_{i}$, $a_{I}=\sum_{i\in I}a_{i}$, and $\rho\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+2}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}$ is the boundary map that glues the last marked points on both components into a node and maps the first $(n_{1}+1)$ points on the first component to the points with the indices $j\not\in I$ and $(n+1)$ and the first $n_{2}$ points on the second component to $i\in I$. ∎
@@
(3.18) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\prod_{i=1}^{m}\psi_{n+i}^{b_{1}}\Xi^{m}_{g,n}|_{u=1}$

### v2

@@ 2.6. The $A=B$ identity in the Gorenstein quotient
As a corollary of this theorem, several important consequences for integrable hierarchies follow. These consequences are developed in detail in [17, Section 4], but we include here a summary of the key ideas for clarity.
To illustrate these implications, consider the case of a cohomological field theory $\{c_{g,n}\}$. The key observation is that the $m$-th derivative of the potential of the CohFT
(2.14) $\frac{\partial^{m}F}{\prod_{i=1}^{m}\partial t^{\alpha_{i},p_{i}}}\big{|}_{t^{1,0}\to t^{1,0}+x}$
is represented as a difference of a naturally defined polynomial in the variables $\{w^{\gamma,d}\}$, introduced in Eq. (1.3), and a series in the variables $\{t^{\gamma,d}+\delta^{\gamma,d}_{1,0}x\}$ whose coefficients are the intersection numbers of the type
(2.15) ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}c_{g,n+m}\big{(}\bigotimes_{i=1}^{n}e_{\beta_{i}}\otimes\bigotimes_{j=1}^{m}e_{\alpha_{i}}\big{)}\prod_{j=1}^{m}\psi_{n+j}^{p_{j}}\mathrm{Coeff}_{a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}}(B^{m}_{g,n}),$
for arbitrary $1\leq\beta_{1},\dots,\beta_{n}\leq N$ and arbitrary $d_{1},\dots,d_{n}\geq 0$ such that $\sum_{i=1}^{n}d_{i}>2g-2+m$. We refer to [17, Theorem 4.6] for a more general version of this observation that works for any F-CohFT, and it is based on [7, Proposition 7.2] and [9, Proposition 3.5].
The first statement of the main theorem implies the existence of the Dubrovin–Zhang hierarchy for any F-CohFT whose classes lie in the tautological ring, in the sense that the system of conservation laws is polynomial [17, Theorem 4.7]. In the special case of a CohFT whose classes belong to the tautological, the vanishing of (2.15) for $m=2$ directly implies the polynomiality of the conservation laws.
The second statement implies that the Dubrovin-Zhang hierarchy is Miura equivalent, in a very explicit way, to the DR hierarchy of conservation laws associated to the same F-CohFT, also under assumption that its classes belong to the tautological ring [17, Theorem 4.10]. In the vein of the discussion above, one has to notice that the integrals
(2.16) ¯$\int_{{\overline{\mathcal{M}}}_{g,n+1}}c_{g,n+1}\big{(}\bigotimes_{i=1}^{n}e_{\beta_{i}}\otimes e_{\alpha}\big{)}\mathrm{Coeff}_{a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}}(A^{1}_{g,n})$
describe the coefficient of a vector field that generates Buryak’s DR hierarchy [17, Theorem 4.9], and thus the corresponding vanishing statement identifies the conservation laws of the the DR hierarchy with the ones of the DZ hierarchy, up to a polynomial correction in the variables $\{w^{\gamma,d}\}$. The $\partial_{x}$ derivative of the latter correction defines the Miura transformation of the two hierarchies [17, Theorem 4.10 and Equation (4.12)] (in the latter reference it is explained in the natural generality of F-CohFTs).
Finally, the third statement implies the strong DR/DZ correspondence for (partial) CohFTs, also under assumption that the classes of (partial) CohFT belong to the tautological ring. In the vein of the discussion above, one has to notice that the integrals
(2.17) ¯$\int_{{\overline{\mathcal{M}}}_{g,n}}c_{g,n}\big{(}\bigotimes_{i=1}^{n}e_{\beta_{i}}\otimes e_{\alpha}\big{)}\mathrm{Coeff}_{a_{1}^{d_{1}}\cdots a_{n}^{d_{n}}}\left(\frac{1}{\sum_{i=1}^{n}a_{i}}\pi_{*}A^{1}_{g,n})\right)$
describe the coefficients of the logarithm of the tau function of Buryak’s DR hierarchy [7, Proposition 6.10] and [8, Theorem 6.1]. This is a consequence of the new expression for the $A$-class, with its correspondence to the original definition detailed in Lemma 2.7. Thus the corresponding vanishing statement identifies logarithms of the tau functions of the DZ and DR hierarchies, up to a polynomial correction in the variables $\{w^{\gamma,d}\}$. The derivatives of the latter correction define the normal Miura transformation that connects the two hierarchies [17, Section 4.4.4] (in the latter reference it is explained in the natural generality of partial CohFTs).
We refer for further details to [17, 7, 9].
@@ 2.7. The trivial CohFT
The second and the third statement are literally equivalent to the statements of the strong DR/DZ correspondence in this case, see [6, 7, 17]. The first statement is equivalent by [17, Section 4] to the fact that the second and higher derivatives of the potential $F(t_{0},t_{1},\dots)$ corresponding to the string solution are known to be differential polynomials in the dependent variables in this case, see e. g. [19]. ∎
@@ 3.1.2. Classes assigned to vertices
where the negative cohomological degree and the space $R^{-1}({\overline{\mathcal{M}}}_{0,2})$ are just formally assigned to allow to treat this case non-exceptionally in what follows, where it is replaced by a contraction rule under the boundary pushforward map. For instance, in this vein it is often convenient to formally extend the definition of the integrals of $\psi$ classes over ¯${\overline{\mathcal{M}}}_{0,2}$, where they are defined by ¯$\int_{{\overline{\mathcal{M}}}_{0,2}}\big{(}(1-x_{1}\psi_{1})(1-x_{2}\psi_{2})\big{)}^{-1}\coloneqq(x_{1}+x_{2})^{-1}$
For a non-root vertex $v$, we consider the class
@@
Note that in this case there exactly $|H(v)|=|H_{+}(v)|+1$ half-edges attached to $v$ with all positive half-edges being the regular legs, and the only negative half-edge being a part of the edge connecting $v$ to $v_{r}$. Note also that in this case $a(v)=a(h_{1})+\cdots+a(h_{|H_{+}(v)|})$. In the exceptional unstable case $g(v)=0,H_{+}(v)=1$ we formally assign to this vertex the following class:
@@
where, as it was for the root vertex, the negative cohomological degree and the space $R^{-1}({\overline{\mathcal{M}}}_{0,2})$ are just formally assigned to allow to treat this case non-exceptionally in what follows.
@@ 3.1.3. Classes assigned to trees
(3.5) $\Xi(T)\coloneqq u^{2g-2+m}\left(\prod_{e\in E(T)}\frac{a(e)}{u}\right)(b_{T})_{*}\left(\left(\sum_{d=-1}^{\infty}\frac{\Psi(v_{r})_{d}}{(-u)^{d}}\right)\otimes\bigotimes_{v\in V(T)\setminus\{v_{r}\}}\left(\sum_{d=-1}^{\infty}\frac{{\mathbb{D}}(v)_{d}}{u^{d}}\right)\right),$
where $(b_{T})_{*}$ is the boundary pushforward map from $\bigotimes_{v\in V(T)}R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})$ to $R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ extended by linearity to the rational functions in $a_{i}$ and Laurent polynomials in $u$. In the case of pre-stable trees $b_{T}$ is assumed to contract the unstable components. Note that under the assumption $02g-2+n+m>$ 0 the dependence of the resulting formula on $a_{i}$’s is purely polynomial, so $\Xi(T)\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q[u,u^{-1}]$. Note also that for $02g-2+n+m>$ 0 the dimension of ¯${\overline{\mathcal{M}}}_{g,n+m}$ is $3g-3+n+m$ and thus the possible range of degrees in $u$ is from $-g+1-n$ to $2g-2+m$.
@@
Note that $\mathrm{Coeff}_{u^{2g-2+m-d}}\Xi(T)\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $\Xi(T)_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$. (Here and below $\mathrm{Coeff}_{u^{p}}\xi$ denotes the coefficient of $u^{p}$ in a Laurent polynomial $\xi(u)$.)
@@ 3.1.4. Classes assigned to sets of trees
Note that $(\Xi^{m}_{g,n})_{d}\coloneqq\mathrm{Coeff}_{u^{2g-2+m-d}}\Xi^{m}_{g,n}\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $(\Xi^{m}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
@@ 3.2. Intersection with a divisor
The other type of the boundary maps $\rho\colon{\overline{\mathcal{M}}}_{g_{1},k_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},k_{2}+1}\to{\overline{\mathcal{M}}}_{g_{1}+g_{2},n+m}$ is indexed by the parameters $g_{1},g_{2}\geq 0$ such that $g_{1}+g_{2}=g$ and the sets of labels $I_{1}\sqcup I_{2}=\{1,\dots,n+m\}$ such that $|I_{i}|=k_{i}$ and $02g_{i}-1+k_{i}>$ 0 for $i=1,2$. We always assume that the marked point labeled by $k_{i}+1$ on ¯${\overline{\mathcal{M}}}_{g_{i},k_{i}+1}$ is the node, and the other $k_{i}$ points are ordered in such a way that the one-to-one map $\{1,\dots,k_{i}\}\to I_{i}\subset\{1,\dots,n+m\}$ preserves the order.
Let $\rho_{2}\colon{\overline{\mathcal{M}}}_{g_{1},k_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},k_{2}+1}\to{\overline{\mathcal{M}}}_{g_{1}+g_{2},n+m}$ be the boundary map such that $m_{1}=|I_{1}\cap\{n+1,\dots,m\}|$ and $m_{2}=|I_{2}\cap\{n+1,\dots,m\}|$ are both non-zero for the corresponding sets of labels $I_{1}$ and $I_{2}$. Let $n_{1}=|I_{1}\cap\{1,\dots,n\}|$ and $n_{2}=|I_{2}\cap\{1,\dots,n\}|$.
@@
Let $\rho_{3}\colon{\overline{\mathcal{M}}}_{g_{1},k_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},k_{2}+1}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map and assume that all points with the labels ${n+1},\dots,{n+m}$ are on the first component, that is, $I_{1}\supset\{n+1,\dots,m\}$ and $I_{1}\cap\{n+1,\dots,m\}=\emptyset$. As before, let $n_{1}=|I_{1}\cap\{1,\dots,n\}|$ and $n_{2}=|I_{2}|\subset\{1,\dots,n\}$.
@@
$\quad+\sigma_{*}(\Xi^{m}_{g_{1},n_{1}+1})_{d-2g_{2}}(\{a_{j}\}_{j\not\in I_{2}},a_{I_{2}})\otimes\lambda_{g_{2}}\mathrm{DR}_{g_{2}}\Big{(}\{a_{i}\}_{i\in I_{2}},-a_{I_{2}}\Big{)}.$
In this formula we assume that the arguments of $\Xi$-classes correspond to the marked points on the corresponding components. In the last summand we use $a_{I_{2}}\coloneqq\sum\nolimits_{i\in I_{2}}a_{i}$. The map $\sigma_{*}$ is induced by the relabeling of the marked points
(3.14) $(1,\dots,n_{1}+m+1)\mapsto(1,\dots,n_{1},n_{1}+m+1,n_{1}+1,\dots,n_{1}+m)$
on ¯${\overline{\mathcal{M}}}_{g_{1},k_{1}+1}$.
@@ 3.3. Intersection with a kappa class
Lemmata in Section 3.2 reduce the vanishing of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha(\Xi^{m}_{g,n})_{d}$ for $d\geq 2g-1+m$ for an arbitrary tautological class $\alpha$ to the vanishing of the same intersection numbers for the classes $\kappa_{c_{1},\dots,c_{l}}\prod_{i=1}^{n+m}\psi_{i}^{b_{i}}$. Indeed, any tautological class $\alpha$ is given as a sum of classes $(b_{\Gamma})_{*}(\beta)$, where $\Gamma$ is a stable graph, $(b_{\Gamma})_{*}$ is the corresponding boundary pushforward map, and $\beta$ is a tensor product of the products of $\psi$- and $\kappa$-classes assigned to the vertices of $\Gamma$. Using the projection formula we can rewrite ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}(b_{\Gamma})_{*}(\beta)(\Xi^{m}_{g,n})_{d}$ as the product over the moduli spaces corresponding to the vertices of $\Gamma$ of the integrals of the classes obtained as components of $(b_{\Gamma})^{*}(\Xi^{m}_{g,n})_{d}$, whose computation then follows the splitting formulas established in Lemmata in Section 3.2, multiplied for the products of $\psi$- and $\kappa$-classes. The latter Lemmata imply that $(b_{\Gamma})^{*}(\Xi^{m}_{g,n})_{d}$ is an external tensor product of the classes $(\Xi^{m^{\prime}}_{g^{\prime},n^{\prime}})_{d^{\prime}}$ and $\lambda_{g}^{\prime}\mathrm{DR}_{g^{\prime}}$. Moreover, for at least one vertex we do obtain as the factor in the external tensor produc the class $(\Xi^{m^{\prime}}_{g^{\prime},n^{\prime}})_{d^{\prime}}$ with the condition $d^{\prime}\geq 2g^{\prime}-1+m^{\prime}$. Thus, once we know that all integrals ¯$\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+m^{\prime}}}\kappa_{c_{1},\dots,c_{l}}\prod_{i=1}^{n^{\prime}+m^{\prime}}\psi_{i}^{b_{i}}(\Xi^{m^{\prime}}_{g^{\prime},n^{\prime}})_{d}$ for $d^{\prime}\geq 2g^{\prime}-1+m^{\prime}$ vanish, we obtain the vanishing of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha(\Xi^{m}_{g,n})_{d}$ for $d\geq 2g-1+m$ for an arbitrary tautological class $\alpha$.
To this end, we have the following statement:
@@ 3.4. Intersection with a $\psi$ class at a regular leg
where $a=\sum_{i=1}^{n}a_{i}$, $a_{I}=\sum_{i\in I}a_{i}$, and $\rho\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+2}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}$ is the boundary map that glues the last marked points on both components into a node and maps the first $(n_{1}+1)$ points on the first component to the points with the indices $j\not\in I$ and $(n+1)$ and the first $n_{2}$ points on the second component to $i\in I$.
In order to see how  (3.16) produces all terms on the right hand side of (3.16), let us focus on a particular graph in $\mathrm{PSSRT}_{g,n,m}$ in the formula for $(\Xi^{m}_{g,n})_{d}$ and fix all degrees of $\psi$-classes associated via $\Psi(v_{r})$ and ${\mathbb{D}}(v)$, $v\in V(T)\setminus\{v_{r}\}$, to the half-edges in $E(T)$, except for one particular half-edge $e_{1}$ that connects $v_{r}$ to a vertex $v_{1}$. For simplicity assume both $v_{r}$ and $v_{1}$ are stable vertices, and the edge $e_{1}$ consists of two half-edges, $h_{1}$ attached to $v_{r}$ and $h_{2}$ attached to $v_{1}$, respectively. Since all degrees in the rest of the graph are fixed and $d$ is fixed, this edge is decorated by $a(e_{1})^{p}\sum_{j=0}^{p}(-1)^{j}\psi(h_{1})^{j}\psi(h_{2})^{p-j}$. Assume further that $\sigma_{i}\in DL(e_{1})$, that is, the leg $i$ is attached to $v_{1}$.
Now, apply (3.18) to $a_{i}\psi_{i}\lambda_{g(v_{1})}\mathrm{DR}_{g(v_{1})}$. The term that corresponds to $-a_{n+1}\psi_{n+1}\lambda_{g}\mathrm{DR}_{g}$ in (3.18) turns the class $a(e_{1})^{p}\sum_{j=0}^{p}\psi(h_{1})^{j}\psi(h_{2})^{p-j}$ associated to the edge $e_{1}$ into the sum of $-a(e_{1})^{p+1}\sum_{j=0}^{p+1}(-1)^{j}\psi(h_{1})^{j}\psi(h_{2})^{p+1-j}$ and $a(e_{1})^{p+1}\psi(h_{1})^{p+1}$. The first term gives a summand in the formula for $-(\Xi^{m}_{g,n})_{d+1}$ on the right hand side of (3.16), and the second term gives a summand in the second line of (3.16) for $g_{2}=g(v_{1})$ and $DL(e_{1})=\{\sigma_{k}\}_{k\in I}$.
The cases when either $v_{r}$ or $v_{1}$ is unstable are analyzed in a similar way. In particular, is $v_{1}$ is unstable, then $DL(e_{1})=\{e_{i}\}$, and we shouldn’t use (3.16), but rather contract the unstable vertex $v_{1}$, replace $h_{1}$ with the leg $\sigma_{i}$, and then note that multiplication by $a_{i}\psi_{i}$ increases the degree of the $\psi$-class assigned to $\sigma_{i}$ by $1$. This gives once again a contribution to the summand $-(\Xi^{m}_{g,n})_{d+1}$ on the right hand side of (3.16). Note that this way we never obtain the degree of the $\psi$-class assigned to $\sigma_{i}$ equal to zero, which must be possible in the formula for $-(\Xi^{m}_{g,n})_{d+1}$, thus we need a counter term for this case hidden in the summand $(\sigma_{i})^{*}(\Xi^{m+1}_{g,n-1}(\{a_{j}\}_{j\not=i}))_{d+1}$ in (3.16). ∎
@@
(3.19) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}\Xi^{m}_{g,n}|_{u=1}$